OpenStudy (anonymous):

Yulian works at the zoo feeding the animals. He puts water in the elephant habitat at the beginning of the day. The table shows how much water, in gallons, remains at different times during the day. What is the equation that represents Yulian’s situation? (y – 29) = (x – ) Yulian’s Situation Hours 2 4 6 8 Water remaining 43 36 29 22

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (studygurl14):

Plug in one of the point from the graph and solve for the missing value (I'll assign variable k to it) (43) - 29 = (2) - k 14 = 2 - k Can you do the rest?